# Decentralized Nonconvex Optimization under Heavy-Tailed Noise: Normalization and Optimal Convergence

### v1

@@ 6 Conclusion and Future Work
In this paper, we have proposed GT-NSGDm for solving decentralized nonconvex smooth optimization to address heavy-tailed noise. The key idea is to leverage normalization, together with momentum variance reduction, to combat heavy-tailed noise, and use gradient tracking to handle cross-node heterogeneity and the nonlinearity brought by normalization. Theoretical analyses establish that GT-NSGDm attains optimal convergence rate when the tail index $p$ is known, and a rate that matches the best centralized one when $p$ is unknown. Extensive experiments on nonconvex linear regression and decentralized Transformer training show that GT-NSGDm is robust and efficient under heavy-tailed noise across various topologies, and achieves a speedup in $n$. Future directions include extending the current analysis to other nonlinearities, such as sign and clipping [17], and generalizing GT-NSGDm to handle objective functions under relaxed smoothness conditions [27].
@@ Appendix A Proof of Claim 1
Consider $n$ scalar functions that for each $i\in\mathcal{V}$, $f_{i}(x)=(1/2)(x-a_{i})^{2}$ for some $a_{i}$, and complete graph with $\bm{W}=(1/n)\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. Let $a_{i}=a,\forall i=1,\ldots,n/2$, and $a_{i}=b,\forall i=n/2+1,\ldots,n$, and $b-a>2B+1$. Let $x_{i}^{0}=a+0.5,\forall i\in\mathcal{V}$. Then, $\forall i\in\mathcal{V}$, vanilla normalization reduces to
@@
Note that the proof above can be further extended to the case where the gradient oracle admits almost surely bounded gradient noise. We can use the noise bound to adapt the choices of $a,b,\varepsilon$ such that all signs still get canceled.
@@ B.2 Intermediate Lemmas
We first present some standard useful relations to be used in our analyses.
@@
=\ = $\|\big{(}\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}\bm{x}^{t}\|$
=$\overset{(i)}{=}$ $\alpha\|\sum_{k=0}^{t-1}\big{(}\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}\big{(}\bm{W}\otimes\bm{I}_{d}\big{)}^{t-k}\mathcal{N}(\bm{y}^{k})\|$
@@
$\overset{(ii)}{\leq}$ $\alpha\|\sum_{k=0}^{t-1}\|\bm{W}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\|_{2}^{t-k}\|\mathcal{N}(\bm{y}^{k})\|$
@@
$\overset{(iii)}{\leq}$ $\frac{\alpha\sqrt{n}\lambda}{1-\lambda}.$ (17)
@@
[…]
@@
${=}$ $\big{(}\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}\big{(}\bm{W}\otimes\bm{I}_{d}\big{)}(\bm{y}^{t-1}+\bm{v}^{t}-\bm{v}^{t-1})$
=\ = $\big{(}\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}\bm{y}^{t-1}+\big{(}\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}(\bm{v}^{t}-\bm{v}^{t-1})$
=$\overset{(i)}{=}$ $\big{(}\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}\big{(}\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}\bm{y}^{t-1}+\big{(}\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}(\bm{v}^{t}-\bm{v}^{t-1})$
=$\overset{(ii)}{=}$ $\sum_{k=0}^{t}\big{(}\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}^{t-k+1}(\bm{v}^{t}-\bm{v}^{t-1}),$ (20)
@@
$\bm{v}^{t}-\bm{v}^{t-1}=(\frac{1}{\beta}-1)(\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\bm{v}^{t})=(\frac{1}{\beta}-1)(\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{k})+\nabla F(\bm{x}^{k})-\bm{v}^{t}).$ (21)
@@
We note that the first half of the right hand side above can be addressed by Lemma 5.
$\mathbb{E}\Big{[}\|\sum_{k=0}^{t}\big{(}\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big{)}^{t-k+1}\big{(}\bm{g}(\bm{x}^{k},\bm{\xi}^{k})-\nabla F(\bm{x}^{k})\big{)}\|\mid\mathcal{F}_{t-1}\Big{]}$
[…]
$\overset{(i)}{\leq}$ $2\sqrt{2}\mathbb{E}\Big{[}\Big{(}\sum_{k=0}^{t}\sum_{i=1}^{n}\lambda^{(t-k+1)p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k})\|^{p}\Big{)}^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\Big{]}$
$\overset{(ii)}{\leq}$ $2\sqrt{2}\Big{(}\mathbb{E}\big{[}\sum_{i=1}^{n}\lambda^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\mid\mathcal{F}_{t-1}\big{]}$
$\quad+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\lambda^{(t-k+1)p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}^{k})\|^{p}\Big{)}^{\frac{1}{p}}$
$\overset{(iii)}{\leq}$ $2\sqrt{2}\Big{(}\lambda^{p}n^{p}\sigma^{p}+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\lambda^{(t-k+1)p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k})\|^{p}\Big{)}^{\frac{1}{p}},$
where we used relation $5$ from Lemma 1 in $(i)$, Jensen’s inequality in $(ii)$, and Assumption 3 in $(iii)$. Applying the above arguments recursively from $\mathcal{F}_{t-2}$ to $\mathcal{F}_{0}$, we have
@@
$=\beta(\bar{\bm{v}}^{t-1}-\overline{\nabla}F(\bm{x}^{t-1}))+(1-\beta)\cdot\underbrace{\frac{1}{n}(\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t}))}_{:=\bm{s}^{t}\in\mathbb{R}^{d}}+\beta\cdot\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}(\bm{x}_{i}^{t-1})-\nabla f_{i}(\bm{x}_{i}^{t}))}_{:=\bm{z}^{t}\in\mathbb{R}^{d}}$
@@
$\begin{split}&\mathbb{E}\Big{[}\|\sum_{k=0}^{t}\beta^{t-k}(1-\beta)\bm{s}^{k}\|\mid\mathcal{F}_{t-1}\Big{]}\\ =\ &\frac{1}{n}\mathbb{E}\Big{[}\|\sum_{k=0}^{t}\sum_{i=1}^{n}\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|\mid\mathcal{F}_{t-1}\Big{]}\\ \leq\ &\frac{2\sqrt{2}}{n}\mathbb{E}\Big{[}\Big{(}\sum_{k=0}^{t}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big{)}^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\Big{]}\\ \overset{(i)}{\leq}\ &\frac{2\sqrt{2}}{n}\Big{(}\mathbb{E}\Big{[}\sum_{k=0}^{t}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\mid\mathcal{F}_{t-1}\Big{]}\Big{)}^{\frac{1}{p}}\\ \leq\ &\frac{2\sqrt{2}}{n}\Big{(}\mathbb{E}\Big{[}\sum_{i=1}^{n}(1-\beta)^{p}\|(\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t}))\|^{p}\mid\mathcal{F}_{t-1}\Big{]}\\ &\quad+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big{)}^{\frac{1}{p}}\\ \overset{(ii)}{\leq}\ &\frac{2\sqrt{2}}{n}\Big{(}n(1-\beta)^{p}\sigma^{p}+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big{)}^{\frac{1}{p}},\end{split}$
where we used Jensen’s inequality in $(i)$ and Assumption 3 in $(ii)$. Recursively applying the preceding arguments from $\mathcal{F}_{t-2}$ to $\mathcal{F}_{0}$, we have
@@
$\mathbb{E}\big{[}\|\bm{v}^{t}-\nabla F(\bm{x}^{t})\|\big{]}\leq\beta^{t+1}\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|+2\sqrt{2}\Big{(}\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\Big{)}^{\frac{1}{p}}n\sigma+n\sum_{k=0}^{t}\beta^{t-k+1}\Big{(}\frac{2\alpha\lambda}{1-\lambda}+\alpha\Big{)}L.$
@@
$\bm{v}^{t}-\nabla F(\bm{x}^{t})$ $=\beta(\bm{v}^{t-1}-\nabla F(\bm{x}^{t-1})+(1-\beta)\underbrace{(\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t}))}_{:=\tilde{\bm{s}}^{t}\in\mathbb{R}^{nd}}+\beta\underbrace{(\nabla F(\bm{x}^{t-1})-\nabla F(\bm{x}^{t}))}_{:=\tilde{\bm{z}}^{t}\in\mathbb{R}^{nd}}$

### v2

@@ 6 Conclusion and Future Work
In this paper, we have proposed GT-NSGDm for solving decentralized nonconvex smooth optimization to address heavy-tailed noise. The key idea is to leverage normalization, together with momentum variance reduction, to combat heavy-tailed noise, and use gradient tracking to handle cross-node heterogeneity and the nonlinearity introduced by normalization. Theoretical analyses establish that GT-NSGDm attains an optimal convergence rate when the tail index $p$ is known, and a rate that matches the best centralized one when $p$ is unknown. Extensive experiments on nonconvex linear regression and decentralized Transformer training show that GT-NSGDm is robust and efficient under heavy-tailed noise across various topologies, and achieves a speedup in $n$. Future directions include extending the current analysis to other nonlinearities, such as sign and clipping [17], and generalizing GT-NSGDm to handle objective functions under relaxed smoothness conditions [27].
Acknowledgments
We thank the anonymous reviewers for their valuable feedback that improved the presentation of this paper and for identifying minor errors in the proofs. The work of S. Yu and S. Kar is supported in part by NSF grant ECCS 2330196. The work of D. Jakovetic was supported by the Ministry of Science, Technological Development, and Innovation (Grants No. 451-03-33/2026-03/ 200125 & 451-03-34/2026-03/ 200125); and the Science Fund of Republic of Serbia, project “LASCADO” (grant 7359).
@@ Appendix A Proof of Claim 1
Consider $n$ scalar functions that for each $i\in\mathcal{V}$, $f_{i}(x)=(1/2)(x-a_{i})^{2}$ for some $a_{i}$, and complete graph with $\bm{W}=(1/n)\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. Let $a_{i}=a,\forall i=1,\ldots,n/2$, and $a_{i}=b,\forall i=n/2+1,\ldots,n$, and ba>2B+1b-a>2B+1. Let $x_{i}^{0}=a+0.5,\forall i\in\mathcal{V}$. Then, $\forall i\in\mathcal{V}$, vanilla normalization reduces to
@@
Note that the proof above can be further extended to the case where the gradient oracle admits almost surely bounded gradient noise. We can use the noise bound to adapt the choices of $a,b,\varepsilon$ such that all signs still get canceled. Similar examples have been used to show divergence results in [62].
@@ Appendix B Proofs of Theorems
Proof structure. The central recursion of our analysis leverages the descent lemma for $L$ smooth functions applied to consecutive network averages of $\{\bm{x}_{i}^{t+1}\}$ and $\{\bm{x}_{i}^{t}\}$ (Lemma 3). This recursion involves two coupled error sources: consensus errors between $\{\bm{x}_{i}^{t}\}$, $\{\bm{y}_{i}^{t}\}$, and gradient estimation errors. We establish the intricate coupling between these errors through a series of intermediate lemmas. Our proof strategy proceeds as follows: we first derive two key lemmas that bound consensus errors in terms of gradient estimation errors (Lemma 4 and 6), then decompose gradient estimation errors into constituent components, including average gradient estimation errors (Lemma 7) and stack gradient estimation errors (Lemma 8), and bound each separately. With all error bounds established, we substitute these into the main recursion and optimize hyperparameter selection to achieve the final order-optimal convergence rates.
@@ B.2 Intermediate Lemmas
We first present some standard useful relations to be used in our analysis.
@@
$=\$$\|\big(\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)\bm{x}^{t}\|$
$\overset{(i)}{=}\$$\alpha\|\sum_{k=0}^{t-1}\big(\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)\big(\bm{W}\otimes\bm{I}_{d}\big)^{t-k}\mathcal{N}(\bm{y}^{k})\|$
@@
$\overset{(ii)}{\leq}\$$\alpha\|\sum_{k=0}^{t-1}\|\bm{W}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\|_{2}^{t-k}\|\mathcal{N}(\bm{y}^{k})\|$
@@
$\overset{(iii)}{\leq}\$$\frac{\alpha\sqrt{n}\lambda}{1-\lambda}.$ (17)
@@
$\leq\$$2\sqrt{2}n^{\frac{1}{2}}\big(\frac{1}{\beta}-1\big)\Big(\sum_{k=0}^{t}\lambda^{(t-k+1)p}\Big)^{\frac{1}{p}}\sigma+\frac{1}{\sqrt{n}}\big(\frac{1}{\beta}-1\big)\sum_{k=0}^{t}\lambda^{t-k+1}\mathbb{E}\big[\|\nabla F(\bm{x}^{k})-\bm{v}^{k}\|\big].$
@@
[…]
$=\$$\big(\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)\bm{y}^{t-1}+\big(\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)(\bm{v}^{t}-\bm{v}^{t-1})$
$\overset{(i)}{=}\$$\big(\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)\big(\bm{I}_{nd}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)\bm{y}^{t-1}+\big(\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)(\bm{v}^{t}-\bm{v}^{t-1})$
$\overset{(ii)}{=}\$$\sum_{k=0}^{t}\big(\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)^{t-k+1}(\bm{v}^{t}-\bm{v}^{t-1}),$ (20)
@@
$\bm{v}^{t}-\bm{v}^{t-1}=(\frac{1}{\beta}-1)(\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\bm{v}^{t})=(\frac{1}{\beta}-1)(\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t})+\nabla F(\bm{x}^{t})-\bm{v}^{t}).$ (21)
@@
We note that the first half of the right hand side above can be addressed by Lemma 5:
$\begin{split}&\mathbb{E}\Big[\|\sum_{k=0}^{t}\big(\bm{W}\otimes\bm{I}_{d}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\otimes\bm{I}_{d}\big)^{t-k+1}\big(\bm{g}(\bm{x}^{k},\bm{\xi}^{k})-\nabla F(\bm{x}^{k})\big)\|\Big]\\ \leq\ &2\sqrt{2}\mathbb{E}\Big[\Big(\sum_{k=0}^{t}\lambda^{(t-k+1)p}\|\bm{g}(\bm{x}^{k},\bm{\xi}^{k})-\nabla F(\bm{x}^{k})\|^{p}\Big)^{\frac{1}{p}}\Big].\end{split}$ (23)
We observe that
$\begin{split}&2\sqrt{2}\mathbb{E}\Big[\Big(\sum_{k=0}^{t}\lambda^{(t-k+1)p}\|\bm{g}(\bm{x}^{k},\bm{\xi}^{k})-\nabla F(\bm{x}^{k})\|^{p}\Big)^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\Big]\\ \leq\ &2\sqrt{2}\mathbb{E}\Big[\Big(\sum_{k=0}^{t}\lambda^{(t-k+1)p}\big(\sum_{i=1}^{n}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k})\|\big)^{p}\Big)^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\Big]\\ \overset{(i)}{\leq}\ &2\sqrt{2}\mathbb{E}\Big[\Big(\sum_{k=0}^{t}\sum_{i=1}^{n}\lambda^{(t-k+1)p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k})\|^{p}\Big)^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\Big]\\ \overset{(ii)}{\leq}\ &2\sqrt{2}\Big(\mathbb{E}\big[\sum_{i=1}^{n}\lambda^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\mid\mathcal{F}_{t-1}\big]\\ &\quad+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\lambda^{(t-k+1)p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}^{k})\|^{p}\Big)^{\frac{1}{p}}\\ \overset{(iii)}{\leq}\ &2\sqrt{2}\Big(\lambda^{p}n^{p}\sigma^{p}+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\lambda^{(t-k+1)p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k})\|^{p}\Big)^{\frac{1}{p}},\end{split}$ (24)
where we used relation $5$ from Lemma 1 in $(i)$, Jensen’s inequality in $(ii)$, and Assumption 3 in $(iii)$. From (23), taking expectations on both sides of (24), and applying the above arguments recursively from $\mathcal{F}_{t-2}$ to $\mathcal{F}_{0}$, we have
@@
$=\beta(\bar{\bm{v}}^{t-1}-\overline{\nabla}F(\bm{x}^{t-1}))+(1-\beta)\cdot\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t}))}_{:=\bm{s}^{t}\in\mathbb{R}^{d}}+\beta\cdot\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}(\bm{x}_{i}^{t-1})-\nabla f_{i}(\bm{x}_{i}^{t}))}_{:=\bm{z}^{t}\in\mathbb{R}^{d}}$
@@
$\begin{split}&\mathbb{E}\Big[\|\sum_{k=0}^{t}\beta^{t-k}(1-\beta)\bm{s}^{k}\|\Big]\\ =\ &\frac{1}{n}\mathbb{E}\Big[\|\sum_{k=0}^{t}\sum_{i=1}^{n}\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|\Big]\\ \leq\ &\frac{2\sqrt{2}}{n}\mathbb{E}\Big[\Big(\sum_{k=0}^{t}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big)^{\frac{1}{p}}\Big].\end{split}$ (28)
Note that
$\begin{split}&\frac{2\sqrt{2}}{n}\mathbb{E}\Big[\Big(\sum_{k=0}^{t}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big)^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\Big]\\ \overset{(i)}{\leq}\ &\frac{2\sqrt{2}}{n}\Big(\mathbb{E}\Big[\sum_{k=0}^{t}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\mid\mathcal{F}_{t-1}\Big]\Big)^{\frac{1}{p}}\\ \leq\ &\frac{2\sqrt{2}}{n}\Big(\mathbb{E}\Big[\sum_{i=1}^{n}(1-\beta)^{p}\|(\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t}))\|^{p}\mid\mathcal{F}_{t-1}\Big]\\ &\quad+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big)^{\frac{1}{p}}\\ \overset{(ii)}{\leq}\ &\frac{2\sqrt{2}}{n}\Big(n(1-\beta)^{p}\sigma^{p}+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\|\beta^{t-k}(1-\beta)(\bm{g}_{i}(\bm{x}_{i}^{k},\bm{\xi}_{i}^{k})-\nabla f_{i}(\bm{x}_{i}^{k}))\|^{p}\Big)^{\frac{1}{p}},\end{split}$ (29)
where we used Jensen’s inequality in $(i)$ and Assumption 3 in $(ii)$. From (28), taking expectations on (29), and recursively applying the preceding arguments from $\mathcal{F}_{t-2}$ to $\mathcal{F}_{0}$, we have
@@
$\mathbb{E}\big[\|\bm{v}^{t}-\nabla F(\bm{x}^{t})\|\big]$
[…]
@@
$\bm{v}^{t}-\nabla F(\bm{x}^{t})$ $=\beta(\bm{v}^{t-1}-\nabla F(\bm{x}^{t-1}))+(1-\beta)\underbrace{(\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t}))}_{:=\tilde{\bm{s}}^{t}\in\mathbb{R}^{nd}}+\beta\underbrace{(\nabla F(\bm{x}^{t-1})-\nabla F(\bm{x}^{t}))}_{:=\tilde{\bm{z}}^{t}\in\mathbb{R}^{nd}}$